# Predictive inference for time series: why is split conformal effective despite temporal dependence?

### v1

@@ 1 Introduction
In spite of these developments, the vanilla split conformal algorithm (Papadopoulos et al., 2002; Lei et al., 2018)—without any modifications or constraints on its implementation—remains an appealing choice for uncertainty quantification in time series models because of its low computational cost and effective practical performance (Chernozhukov et al., 2018; Xu and Xie, 2023b; Oliveira et al., 2024). Indeed, the accurate predictive coverage of split conformal prediction on time series data that has often been observed empirically may seem quite surprising: due to temporal dependence, time series data is generally far from exchangeable, so how can a framework whose justification relies on exchangeability perform so well? The purpose of this paper is to explain the (often) strong performance of this algorithm in the time series setting.
@@ 1.1 The predictive inference problem
Split conformal prediction (Papadopoulos et al., 2002; Vovk et al., 2005) addresses this problem with the following method. Suppose we have a score function $s:\mathcal{Z}\to\mathbb{R}$ that we can evaluate on our data points. Assume for the moment that $s$ is pretrained—that is, the definition of $s$ does not depend on $\mathbf{Z}$. Treating our first $n$ data points as calibration data, we observe that if the $n+1$ data points are iid, the score evaluated at the test point, $s(Z_{n+1})$, must conform to the scores of the calibration data points, $(s(Z_{i}))_{i\in[n]}$ (in that it must be drawn from the same distribution). If we wish to guarantee coverage with probability at least $1-\alpha$, the split conformal prediction set is then given by
@@
where the correction factor $1+\nicefrac{{1}}{{n}}$ to the coverage is to account for the fact that we can only compute the quantile on the $n$ training points without including the test point. [^1] To formally define the notation $\mathrm{Quantile}(\cdot)$, which computes the quantile of a finite list of values, for any $v\in\mathbb{R}^{m}$ we use $\mathrm{Quantile}_{b}(v)$ to denote the $\lceil bm\rceil$-th order statistic of the vector, i.e., $v_{(\lceil bm\rceil)}$ where $v_{(1)}\leq\dots\leq v_{(m)}$. We will use the convention that $\mathrm{Quantile}_{b}(v)=\infty$ if $b>1$, and $\mathrm{Quantile}_{b}(\infty)=$-$\infty$ if $b\leq 0$. A canonical example in the setting of a real-valued response ($\mathcal{Y}=\mathbb{R}$) is the regression score, $s(z)=|y-\widehat{f}(x)|$ where $z=(x,y)$ and $\widehat{f}$ is a pretrained regression model. This leads to a prediction set of the form $\widehat{C}_{n}(X_{n+1})=\widehat{f}(X_{n+1})\pm\mathrm{Quantile}_{(1-\alpha)(1+1/n)}(s(Z_{1}),\dots,s(Z_{n}))$. However, the split conformal method may be implemented with any score function.
@@ 1.2 A motivating numerical experiment
To see how conformal prediction can perform well in a time series setting, let us illustrate the coverage attained by the pretrained approach on a toy example. Let $\{W_{j}\}_{j\in\mathbb{Z}}$ denote a collection of standard Gaussian variables, and for each $i\in[n+1]$, set $\epsilon_{i}=\sum_{j=i-t}^{i}W_{j}$ to be a moving average process of order $t$ with unit coefficients; denote the joint distribution of $(\epsilon_{i})_{i\in[n+1]}$ by $\mathsf{MA}(t;\mathbf{1})$. Suppose we have a time series of data $(X_{i},Y_{i})_{i\in[n+1]}$ generated from the standard regression model
@@
Figure 1: Coverage of the pretrained conformal prediction set (1) on a sequence of length $n$ from the moving average process (3) of order $t$. The desired (i.e. nominal) coverage is $90\%$ throughout, and is denoted by a dotted line in all plots. Each point is generated by averaging over $10^{6}$ empirical trials.
In Figure 1, we plot various properties of the coverage achieved by this prediction interval. Clearly, the prediction interval achieves the desired coverage if the MA process has order $t=0$, in which case the process is iid, but for all other settings it suffers from a loss of coverage. Based on these plots, we might conjecture that the loss in coverage for split conformal prediction is proportional to $\nicefrac{{t}}{{n}}$. But can we guarantee that the coverage loss is always bounded in this fashion? This paper will provide an affirmative answer to this question for a larger class of time series models, accommodating not just pretrained scores and memoryless predictors but also the split conformal approach (2) and predictors with memory, which we introduce next.
@@ 1.3 Pretrained and split conformal for predictors with memory
$\begin{split}\widehat{C}_{n}(X_{n+1};&Z_{n},\dots,Z_{n-L+1})={}\\ &\left\{y\in\mathcal{Y}:s((X_{n+1},y);Z_{n},\dots,Z_{n-L+1})\leq\mathrm{Quantile}_{(1-\alpha)(1+\frac{1}{n-L})}(S_{L+1},\dots,S_{n})\right\}\end{split}$ (4a)
@@
$\begin{split}\widehat{C}_{n}(X_{n+1};&Z_{n},\dots,Z_{n-L+1})={}\\ &\left\{y\in\mathcal{Y}:s((X_{n+1},y);Z_{n},\dots,Z_{n-L+1})\leq\mathrm{Quantile}_{(1-\alpha)(1+\frac{1}{n_{1}-L})}(S_{n_{0}+L+1},\dots,S_{n})\right\},\end{split}$ (5)
@@ 1.4 Related work
Moving beyond split conformal, other methods have been specifically developed for the time series setting (and more broadly for non-exchangeable settings). Notable examples are conformal prediction algorithms due to Chernozhukov et al. (2018, 2021), which rely on approximate block exchangeability of time series data and ensemble methods due to Xu and Xie (2023a), which are proven to work when we have a consistent predictor. Other methods are based on weighted versions of conformal prediction (Tibshirani et al., 2019; Fannjiang et al., 2022; Prinster et al., 2024), but these approaches involve correcting for a known distribution shift or temporal dependence—information that is not typically available for most time series data. A final family of methods is derived from online learning (e.g., Gibbs and Candès, 2021, 2024), and views the construction of uncertainty sets as a game between nature and the statistician.
@@ 1.5 Contributions and organization
  • We introduce the notion of a switch coefficient for a dependent stochastic process, which measures the total variation distance when we swap certain subvectors of the time series of data points. We show that the switch coefficients can be bounded for $\beta$-mixing processes—and consequently, processes such as the one in the motivating example (3) are covered by our theory.
  • We bound the coverage loss of pretrained conformal prediction by a function of the switch coefficient of the score process. For the MA process and its relatives, this result theoretically confirms the empirical observation made in Figure 1, and holds over a more general class of stochastic processes while accommodating predictors with memory. Moreover, we show that our characterization is tight over the class of stationary, $\beta$ mixing sequences.
@@ 2.1 The switch coefficients
which is the subvector consisting of the first $m-\tau-k$ entries of $\mathbf{w}$ followed by the last $k$ entries of $\mathbf{w}$, and is obtained by deleting a block of $\tau$ many entries after position $m-\tau-k$. Similarly, we define
@@ 2.2 Connection to mixing coefficients
where $\mathbf{Z}^{\prime}=(Z^{\prime}_{1},\dots,Z^{\prime}_{n+1})\in\mathcal{Z}^{n+1}$ denotes an iid copy of $\mathbf{Z}$. In other words, if $\beta(\tau)$ is small, this means that the subvectors $(Z_{1},\dots,Z_{k})$ and $(Z_{k+\tau+1},\dots,Z_{n+1})$ are approximately independent.
@@ 2.3 Switching data points, or switching scores?
Suppose we are working with a pretrained score function $s$. Since the prediction set $\widehat{C}_{n}$ depends on the data points only through their scores, we may ask whether the time series of scores is approximately exchangeable. How does this question relate to the properties of the data time series $\mathbf{Z}$ itself?
@@
In contrast, in the general case with memory $L\geq 0$, the situation is somewhat more complicated. For example, even if the data points $Z_{i}$ are exchangeable, the scores are not exchangeable when the memory $L$ is positive, and indeed may have strong temporal dependence. In particular, writing $\mathbf{S}=(S_{L+1},\dots,S_{n+1})$ where $S_{i}=s(Z_{i};Z_{i-1},\dots,Z_{i-L})$, we may have $\bar{\Psi}_{\tau}(\mathbf{Z})=0$ but $\bar{\Psi}_{\tau}(\mathbf{S})>0$, unlike in the memoryless case. Nonetheless, we can still relate the switch coefficients of the scores to those of the data:
@@ 3.1 Coverage guarantee for the pretrained setting
Theorem 1 is proved in Section A.1. While Theorem 1 is stated in terms of the switch coefficients of the scores, combining this result with Propositions 1 and 2 immediately yields the following corollary, which characterizes the coverage in terms of the properties of the time series of data points $\mathbf{Z}$.
@@ 3.2 A matching lower bound
This implies that the coverage gap in (6) (and hence, the guarantee given in Theorem 1) is tight up to a factor $\frac{1-\alpha}{4}$. Since it is typical to take $\alpha\leq 1/2$, this factor should be viewed as a universal constant greater than $1/8$.
@@ 3.3 Can the conformal prediction set overcover?
Theorem 3 is proved in Section A.3. We note that as a corollary, upper bounds in terms of the properties of the time series $\mathbf{Z}$ (analogous to Corollary 1) also follow in this case.
@@ 3.4 Coverage guarantee for the split conformal setting
Next, we turn to the setting of split conformal prediction, where the score function $s$ is now trained on a portion of the available data, as in (5). Throughout, we will assume that the sample size is split as $n=n_{0}+n_{1}$, where $n_{0},n_{1}\geq 1$ and $n_{1}\geq L$. Write $\mathbf{S}=(S_{n_{0}+L+1},\dots,S_{n+1})$, the vector of scores on the calibration set together with the test point score, $S_{n+1}=s(Z_{n+1};Z_{n},\dots,Z_{n-L+1})$. Define also
@@
We prove Proposition 3 in Section A.7. As we will see in the proof, the key step is to bound $\Psi_{k,\tau}(\mathbf{S}_{\textnormal{split},\tau_{*}})$ using total variation distances of certain subvectors of $\mathbf{Z}$ (a more complex form of the switch coefficient). Combining this result with Theorem 4, we immediately obtain the following corollary.
@@
Let us compare again with the result of Oliveira et al. (2024, Theorem 4), who show that if the score function is memoryless, then its coverage loss for split conformal prediction on a $\beta$-mixing process is bounded (in our notation and up to logarithmic factors) by a term of the order $\min_{\tau}\{\sqrt{\nicefrac{{\tau}}{{n}}}+\sqrt{\nicefrac{{\tau_{*}}}{{n}}}+2\beta(\tau)+2\beta(\tau_{*})\}$. As before, comparing with Corollary 2 above, note that our bound on the coverage loss is tighter, scaling linearly in $\nicefrac{{\tau}}{{n}}$ and $\nicefrac{{\tau_{*}}}{{n}}$. Once again, this improvement is a consequence of our new proof technique.
@@ 4 Discussion
Motivated by the question of why pretrained and split conformal prediction are effective in spite of temporal dependence, we introduced a new “switch coefficient” to measure the deviation of scores from exchangeability, and showed that the loss of coverage is bounded whenever the score process has small switch coefficient. This covers the class of $\beta$-mixing processes, and improves upon previous characterizations of the coverage loss. We also showed that our characterization of the coverage loss is tight, and can accurately reflect empirically observed behavior in canonical time series models.
@@ A.6 Proof of Proposition 2
Then, by construction, we have $\Delta^{0}_{k,\tau}(\mathbf{S})=f_{k}\big(\Delta^{0}_{k+L,\tau-L}(\mathbf{Z})\big)$ and $\Delta^{1}_{k,\tau}(\mathbf{S})=f_{k}\big(\Delta^{1}_{k+L,\tau-L}(\mathbf{Z})\big)$. Therefore,

### v2

@@ 1 Introduction
In spite of these developments, the vanilla split conformal algorithm (Papadopoulos et al., 2002; Lei et al., 2018)—without any modifications or constraints on its implementation—remains an appealing choice for uncertainty quantification in time series models because of its low computational cost and effective practical performance (Chernozhukov et al., 2018; Xu and Xie, 2023b; Oliveira et al., 2024). On the face of it, this may seem quite surprising: due to temporal dependence, time series data is generally far from exchangeable, so how can a framework whose justification relies on exchangeability perform so well? The purpose of this paper is to explain the (often) strong performance of this algorithm in the time series setting.
@@ 1.1 The predictive inference problem
Split conformal prediction (Papadopoulos et al., 2002; Vovk et al., 2005) addresses this problem with the following method. Suppose we have a score function $s:\mathcal{Z}\to\mathbb{R}$ that we can evaluate on our data points. Assume for the moment that $s$ is pretrained—that is, the definition of $s$ does not depend on $\mathbf{Z}$. Treating our first $n$ data points as calibration data, we observe that if all data points are iid, the score evaluated at the test point, $s(Z_{n+1})$, must conform to the scores of the calibration data points, $(s(Z_{i}))_{i\in[n]}$ (in that it must be drawn from the same distribution). If we wish to guarantee coverage with probability at least $1-\alpha$, the split conformal prediction set is then given by
@@
where the correction factor $1+\nicefrac{{1}}{{n}}$ to the coverage is to account for the fact that we can only compute the quantile on the $n$ training points without including the test point. [^1] To formally define the notation $\mathrm{Quantile}(\cdot)$, which computes the quantile of a finite list of values, for any $v\in\mathbb{R}^{m}$ we use $\mathrm{Quantile}_{b}(v)$ to denote the $\lceil bm\rceil$-th order statistic of the vector, i.e., $v_{(\lceil bm\rceil)}$ where $v_{(1)}\leq\dots\leq v_{(m)}$. We will use the convention that $\mathrm{Quantile}_{b}(v)=\infty$ if $b>1$, and $\mathrm{Quantile}_{b}(v)=$-$\infty$ if $b\leq 0$. A canonical example in the setting of a real-valued response ($\mathcal{Y}=\mathbb{R}$) is the regression score, $s(z)=|y-\widehat{f}(x)|$ where $z=(x,y)$ and $\widehat{f}$ is a pretrained regression model. This leads to a prediction set of the form $\widehat{C}_{n}(X_{n+1})=\widehat{f}(X_{n+1})\pm\mathrm{Quantile}_{(1-\alpha)(1+1/n)}(s(Z_{1}),\dots,s(Z_{n}))$. However, the split conformal method may be implemented with any score function.
@@ 1.2 A motivating numerical experiment
Despite the theoretical requirement of exchangeability, split conformal prediction has often been observed to perform well on time series data, where the standard exchangeability assumptions are strongly violated. This phenomenon has been observed repeatedly in the literature, where split conformal is often observed to be competitive with other uncertainty quantification methods for time series (see, e.g. Oliveira et al., 2024; Xu and Xie, 2023a).
For concreteness, let us illustrate the coverage of split conformal on a toy example. Let $(W_{j})_{j\in\mathbb{Z}}$ denote a collection of standard Gaussian variables, and for each $i\in[n+1]$, set $\epsilon_{i}=\sum_{j=i-t}^{i}W_{j}$ to be a moving average process of order $t$ with unit coefficients; denote the joint distribution of $(\epsilon_{i})_{i\in[n+1]}$ by $\mathsf{MA}(t;\mathbf{1})$. Suppose we have a time series of data $(X_{i},Y_{i})_{i\in[n+1]}$ generated from the standard regression model
@@
Figure 1: Coverage of the pretrained conformal prediction set (1) on a sequence of length $n$ from the moving average process (3) of order $t$. The desired (i.e. nominal) coverage is $90\%$ in both experiments, and is denoted by a dotted line. Each point is generated by averaging over $10^{6}$ empirical trials.
In Figure 1, we plot the coverage achieved by this prediction interval. Clearly, the prediction interval achieves the desired coverage if the MA process has order $t=0$, in which case the process is iid, but for all other settings it suffers from a modest loss of coverage. Based on these plots, we might conjecture that the loss in coverage for split conformal prediction is proportional to $\nicefrac{{t}}{{n}}$. But can we guarantee that the coverage loss is always bounded in this fashion? This paper will provide an affirmative answer to this question for a larger class of time series models, accommodating not just pretrained scores and memoryless predictors but also the split conformal approach (2) and predictors with memory, which we introduce next.
@@ 1.3 Pretrained and split conformal for predictors with memory
$\begin{split}\widehat{C}_{n}(&X_{n+1};Z_{n},\dots,Z_{n-L+1})={}\\ &\left\{y\in\mathcal{Y}:s((X_{n+1},y);Z_{n},\dots,Z_{n-L+1})\leq\mathrm{Quantile}_{(1-\alpha)(1+\frac{1}{n-L})}(S_{L+1},\dots,S_{n})\right\}\end{split}$ (4a)
@@
$\begin{split}&\widehat{C}_{n}(X_{n+1};Z_{n},\dots,Z_{n-L+1})={}\\ &\left\{y\in\mathcal{Y}:s((X_{n+1},y);Z_{n},\dots,Z_{n-L+1})\leq\mathrm{Quantile}_{(1-\alpha)(1+\frac{1}{n_{1}-L})}(S_{n_{0}+L+1},\dots,S_{n})\right\},\end{split}$ (5)
@@ 1.4 Related work
Moving beyond split conformal, other methods have been specifically developed for the time series setting (and more broadly for non-exchangeable settings). Notable examples are conformal prediction algorithms due to Chernozhukov et al. (2018, 2021), which rely on approximate block exchangeability of time series data and ensemble methods due to Xu and Xie (2023a), which are proven to work when we have a consistent predictor. Other methods are based on weighted versions of conformal prediction (Tibshirani et al., 2019; Fannjiang et al., 2022; Prinster et al., 2024), but these approaches involve correcting for a known distribution shift—information that is not typically available for most time series data. A final family of methods is derived from online learning (e.g., Gibbs and Candès, 2021, 2024), and views the construction of uncertainty sets as a game between nature and the statistician.
@@ 1.5 Contributions and organization
  • We introduce the notion of a switch coefficient for a dependent stochastic process, which measures the total variation distance when we swap certain subvectors of the time series. We show that the switch coefficients can be bounded for $\beta$-mixing processes—and consequently, processes such as the one in the motivating example (3) are covered by our theory.
  • We bound the coverage loss of pretrained conformal prediction by a function of the switch coefficient of the score process. For the MA process and its relatives, this result theoretically confirms the empirical observation made in Figure 1, and holds over a more general class of stochastic processes while accommodating predictors with memory. Moreover, we show that our characterization is tight over the class of stationary, $\beta$-mixing sequences.
@@ 2.1 The switch coefficients
which is the subvector consisting of the first $m-\tau-k$ entries of $\mathbf{w}$ followed by the last $k$ entries of $\mathbf{w}$, and is obtained by deleting a block of $\tau$ many entries after position $m-\tau-k$. Similarly, define
@@ 2.2 Connection to mixing coefficients
where $\mathbf{Z}^{\prime}=(Z^{\prime}_{1},\dots,Z^{\prime}_{n+1})\in\mathcal{Z}^{n+1}$ denotes an iid copy of $\mathbf{Z}$. In other words, if $\beta(\tau)$ is small, this means that the subvectors $(Z_{1},\dots,Z_{k})$ and $(Z_{k+\tau+1},\dots,Z_{n+1})$ are approximately independent. Note that the deletion operation (Definition 1), is closely connected to $\beta$-mixing: the $\beta$-mixing condition is defined in terms of measuring dependence between $(Z_{1},\dots,Z_{k})$ and $(Z_{k+\tau+1},\dots,Z_{n+1})$, which involves deleting $\tau$ consecutive data points in the time series.
Next we relate $\beta$-mixing coefficients to the switch coefficients defined above.
@@ 2.3 Switching data points, or switching scores?
Suppose we are working with a pretrained score function $s$. Since the prediction set $\widehat{C}_{n}$ depends on the data points only through their scores, we may ask whether the time series of scores is approximately exchangeable. How does this question relate to the properties of the data time series $\mathbf{Z}$?
@@
In contrast, in the general case with memory $L\geq 0$, the situation is somewhat more complicated. For example, even if the data points $Z_{i}$ are exchangeable, the scores are not exchangeable when the memory $L$ is positive, and indeed may have strong temporal dependence. In particular, writing $\mathbf{S}=(S_{L+1},\dots,S_{n+1})$ where $S_{i}=s(Z_{i};Z_{i-1},\dots,Z_{i-L})$, we may have $\bar{\Psi}_{\tau}(\mathbf{Z})=0$ but $\bar{\Psi}_{\tau}(\mathbf{S})>0$, unlike in the memoryless case. Nonetheless, we can still relate the switch coefficients of the scores to those of the data, as shown in the following proposition.
@@ 3.1 Coverage guarantee for the pretrained setting
Theorem 1 is proved in Section A.1. While Theorem 1 is stated in terms of the switch coefficients of the scores, combining this result with Propositions 1 and 2 immediately yields the following corollary, which characterizes the coverage in terms of the properties of the time series $\mathbf{Z}$.
@@ 3.2 A matching lower bound
This implies that the coverage gap in (6) (and hence, the guarantee given in Theorem 1) is tight up to a factor $\frac{1-\alpha}{4}$. Since it is typical to take $\alpha\leq 1/2$, this factor should be viewed as a universal constant in the range $\nicefrac{{1}}{{8}},1]$.
@@ 3.3 Can the conformal prediction set overcover?
Theorem 3 is proved in Section A.3. We note that as a corollary, upper bounds in terms of the properties of the time series $\mathbf{Z}$ (analogous to Corollary 1) also follow in this case. While the coverage upper bound in Theorem 3 may be conservative for certain settings, it is possible—analogously to Theorem 2—to construct an example that essentially achieves this upper bound.
@@ 3.4 Coverage guarantee for the split conformal setting
Next, we turn to split conformal prediction, where the score function $s$ is now trained on a portion of the available data, as in (5). Throughout, we will assume that the sample size is split as $n=n_{0}+n_{1}$, where $n_{0},n_{1}\geq 1$ and $n_{1}\geq L$. Write $\mathbf{S}=(S_{n_{0}+L+1},\dots,S_{n+1})$, the vector of scores on the calibration set together with the test point score, $S_{n+1}=s(Z_{n+1};Z_{n},\dots,Z_{n-L+1})$. Define also
@@
We prove Proposition 3 in Section A.7. As we will see in the proof, the key step is to bound $\Psi_{k,\tau}(\mathbf{S}_{\textnormal{split},\tau_{*}})$ using total variation distances of certain subvectors of $\mathbf{Z}$ (a more complex form of the switch coefficient). Combining this result with Theorem 4, we obtain the following corollary.
@@
Let us compare again with the result of Oliveira et al. (2024, Theorem 4), who show that if the score function is memoryless, then its coverage loss for split conformal prediction on a $\beta$-mixing process is bounded (in our notation and up to logarithmic factors) by a term of the order $\min_{\tau,\tau_{*}}\{\sqrt{\nicefrac{{\tau}}{{n}}}+\sqrt{\nicefrac{{\tau_{*}}}{{n}}}+2\beta(\tau)+2\beta(\tau_{*})\}$. As before, comparing with Corollary 2 above, note that our bound on the coverage loss is tighter, scaling linearly in $\nicefrac{{\tau}}{{n}}$ and $\nicefrac{{\tau_{*}}}{{n}}$. Once again, this improvement is a consequence of our new proof technique.
@@ 4 Discussion
Motivated by the question of why pretrained and split conformal prediction are effective in spite of temporal dependence, we introduced a new “switch coefficient” to measure the deviation of scores from exchangeability, and showed that the loss of coverage is bounded whenever the score process has small switch coefficient. This covers the class of $\beta$-mixing processes, and improves upon previous characterizations of the coverage loss. We also showed that our characterization of the coverage loss is tight, and can accurately reflect empirically observed behavior in canonical time series models. An important open question is whether these bounds on under- and over-coverage are optimal for any predictive method—that is, whether split conformal is optimal for this class of problems, or whether alternative methods (that may use knowledge of the mixing properties of the time series) could improve these bounds.
@@ A.6 Proof of Proposition 2
Then, by construction, $\Delta^{0}_{k,\tau}(\mathbf{S})=f_{k}\big(\Delta^{0}_{k+L,\tau-L}(\mathbf{Z})\big)$ and $\Delta^{1}_{k,\tau}(\mathbf{S})=f_{k}\big(\Delta^{1}_{k+L,\tau-L}(\mathbf{Z})\big)$. Therefore,